# Integrative Learning of Intensity Fluctuations of Quantum Dots under Excitation via a Tailored Mixture Hidden Markov Model

### v1

@@ A.5 Variance Estimation
Combining the above two parts could give the estimate of the variance-covariance matrix, $\boldsymbol{V}$. Our limited numerical expeiment suggests that this method works properly.
Appendix B Additional Results in Data Analysis
B.1 Results from 3-Cluster Model
@@
B.2 Results from 4-Cluster Model
@@
$\begin{split}\hat{\boldsymbol{\pi}}_{1}&=\begin{bmatrix}0.234&0.453&0.313\end{bmatrix},\quad\hat{\boldsymbol{\pi}}_{2}=\begin{bmatrix}0.000&0.091&0.909\end{bmatrix},\\ \hat{\boldsymbol{\pi}}_{3}&=\begin{bmatrix}0.066&0.342&0.592\end{bmatrix},\quad\hat{\boldsymbol{\pi}}_{4}=\begin{bmatrix}0.155&0.574&0.271\end{bmatrix}.\end{split}$
@@
From the 3-cluster model to the 4-cluster model, the correspondence of the cluster assignments are as follows.
@@
In the 4-cluster model, Clusters 1 and 2 align well with Clusters 1 and 2 in the 3-cluster model. The estimated transition matrices and the equilibrium distributions for these two clusters closely match those from the 3-cluster model. It is clear that Clusters 3 and 4 in the 4-cluster model are mainly from splitting Cluster 3 in the 3-cluster model. Specifically, the new Cluster 3 is associated with a greater proportion of time spent in the high state, while the new Cluster 4 exhibits longer durations in the medium state. From the perspective of the transition matrices, the new Cluster 3 demonstrates a higher probability of transitioning to and remaining in the high state, as evidenced by the last column of $\widehat{\boldsymbol{\Pi}}_{3}$. Conversely, the new Cluster 4 is characterized by a stronger tendency to remain in the medium state compared to the new Cluster 3.
B.3 Results from 5-Cluster Model
$\widehat{\boldsymbol{\Pi}}_{1}=\begin{bmatrix}0.814&0.16&0.026\\ 0.073&0.774&0.153\\ 0.029&0.28&0.692\end{bmatrix},\quad\widehat{\boldsymbol{\Pi}}_{2}=\begin{bmatrix}0.000&1.000&0.000\\ 0.000&0.993&0.007\\ 0.000&0.000&1.000\end{bmatrix},\quad\widehat{\boldsymbol{\Pi}}_{3}=\begin{bmatrix}0.889&0.092&0.019\\ 0.036&0.785&0.179\\ 0.006&0.132&0.863\end{bmatrix},$
@@ B.3 Results from 5-Cluster Model
$\begin{split}&\hat{\boldsymbol{\pi}}_{1}=\begin{bmatrix}0.236&0.497&0.267\end{bmatrix},\quad\hat{\boldsymbol{\pi}}_{2}=\begin{bmatrix}0.000&0.049&0.951\end{bmatrix},\quad\hat{\boldsymbol{\pi}}_{3}=\begin{bmatrix}0.144&0.363&0.493\end{bmatrix},\\[10.0pt] &\hat{\boldsymbol{\pi}}_{4}=\begin{bmatrix}0.219&0.577&0.203\end{bmatrix},\quad\hat{\boldsymbol{\pi}}_{5}=\begin{bmatrix}0.025&0.427&0.548\end{bmatrix}.\end{split}$
@@
From the 4-cluster model to the 5-cluster model, the correspondence of the cluster assignments are as follows.

### v2

@@ B.1 Variance Estimation
Combining two key parts could give the estimate of the variance-covariance matrix, $\boldsymbol{V}$.
Appendix C Additional Simulation Results
C.1 Additional Results from Simulation
We consider a third scenario that is slightly modified from Scenario 1, to make the three clusters slightly more distinct. The transition matrices are given by
$\boldsymbol{\Pi}_{1}=\begin{bmatrix}0.848&0.127&0.025\\ 0.099&0.735&0.166\\ 0.017&0.186&0.796\end{bmatrix},\boldsymbol{\Pi}_{2}=\begin{bmatrix}0.795&0.205&0.000\\ 0.000&0.994&0.006\\ 0.000&0.001&0.999\end{bmatrix},\boldsymbol{\Pi}_{3}=\begin{bmatrix}0.938&0.057&0.005\\ 0.013&0.924&0.063\\ 0.001&0.067&0.932\end{bmatrix},$
and the parameters for the three states are set as
$(a_{1},b_{1},\epsilon_{1}^{(0)},\epsilon_{1}^{(1)})$ $=(2.000,5.000,0.025,0.000),$
$(a_{2},b_{2},\epsilon_{2}^{(0)},\epsilon_{2}^{(1)})$ $=(10.000,7.000,0.000,0.001),$
$(a_{3},b_{3},\epsilon_{3}^{(0)},\epsilon_{3}^{(1)})$ $=(12.000,3.000,0.000,0.020).$
The simulation results are showing in Tables 7 and 8. The observations from these results are all consistent with those reported in Section 4 of the main manuscript; we thus omit the details here.
Table 7: Simulation: Results for Scenario 3 with balanced partition. For better presentation, values in $\mbox{Er}(\widehat{\boldsymbol{\delta}})$ and $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ are multiplied by $10$, and values in $\mbox{Er}(\widehat{\mu})$ and $\mbox{Er}(\widehat{\sigma}^{2})$ are multiplied by $10^{2}$.
Method Length ($T$) $\mbox{Er}(\widehat{\mu})$ $\mbox{Er}(\widehat{\sigma}^{2})$ $\mbox{Er}(\widehat{\boldsymbol{\delta}})$ $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ $\mbox{Er}(\widehat{\boldsymbol{\Pi}})$ CC Balanced, Sample size ($N$) = 100 MHMM-$\beta^{*}$ 250 0.36 (0.02) 0.66 (0.04) 0.00 (0.00) 2.96 (0.12) 0.12 (0.01) HMM-G 14.50 (0.12) 10.15 (0.07) 3.42 (0.08) 2.18 (0.03) 0.54 (0.00) HMM-$\beta$ 21.77 (0.18) 9.23 (0.09) 3.93 (0.06) 2.21 (0.03) 0.53 (0.00) MHMM-G 0.94 (0.04) 2.94 (0.08) 0.89 (0.12) 0.60 (0.07) 0.93 (0.01) MHMM-$\beta$ 0.37 (0.02) 0.66 (0.04) 0.46 (0.10) 3.06 (0.12) 0.30 (0.05) 0.97 (0.01) MHMM-$\beta^{*}$ 500 0.25 (0.01) 0.51 (0.03) 0.00 (0.00) 2.15 (0.09) 0.10 (0.01) HMM-G 14.00 (0.09) 8.68 (0.05) 3.75 (0.05) 2.14 (0.02) 0.51 (0.00) HMM-$\beta$ 21.07 (0.13) 6.24 (0.06) 3.76 (0.05) 1.98 (0.02) 0.52 (0.00) MHMM-G 0.92 (0.03) 2.90 (0.07) 0.93 (0.14) 0.56 (0.06) 0.93 (0.01) MHMM-$\beta$ 0.26 (0.01) 0.52 (0.03) 0.45 (0.11) 2.29 (0.09) 0.31 (0.05) 0.97 (0.01) MHMM-$\beta^{*}$ 1000 0.18 (0.01) 0.36 (0.02) 0.00 (0.00) 1.45 (0.05) 0.10 (0.01) HMM-G 13.07 (0.06) 7.75 (0.03) 3.56 (0.04) 2.01 (0.02) 0.58 (0.01) HMM-$\beta$ 19.63 (0.11) 4.62 (0.06) 3.87 (0.02) 1.95 (0.01) 0.58 (0.00) MHMM-G 0.85 (0.02) 2.79 (0.05) 0.71 (0.13) 0.42 (0.05) 0.95 (0.01) MHMM-$\beta$ 0.21 (0.01) 0.41 (0.02) 0.62 (0.13) 1.61 (0.07) 0.38 (0.06) 0.96 (0.01) MHMM-$\beta^{*}$ 2000 0.13 (0.01) 0.26 (0.01) 0.00 (0.00) 1.11 (0.04) 0.09 (0.01) HMM-G 12.23 (0.05) 7.44 (0.02) 2.87 (0.14) 1.77 (0.04) 0.70 (0.02) HMM-$\beta$ 18.49 (0.08) 3.81 (0.03) 3.41 (0.13) 1.84 (0.03) 0.67 (0.01) MHMM-G 0.92 (0.02) 2.90 (0.03) 0.64 (0.12) 0.36 (0.04) 0.95 (0.01) MHMM-$\beta$ 0.16 (0.01) 0.32 (0.02) 0.30 (0.10) 1.42 (0.07) 0.26 (0.04) 0.98 (0.01)
Table 8: Simulation: Results for Scenario 3 with unbalanced partition. For better presentation, values in $\mbox{Er}(\widehat{\boldsymbol{\delta}})$ and $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ are multiplied by $10$, and values in $\mbox{Er}(\widehat{\mu})$ and $\mbox{Er}(\widehat{\sigma}^{2})$ are multiplied by $10^{2}$.
Method Length ($T$) $\mbox{Er}(\widehat{\mu})$ $\mbox{Er}(\widehat{\sigma}^{2})$ $\mbox{Er}(\widehat{\boldsymbol{\delta}})$ $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ $\mbox{Er}(\widehat{\boldsymbol{\Pi}})$ CC Unbalanced, Sample size ($N$) = 100 MHMM-$\beta^{*}$ 250 0.34 (0.02) 0.60 (0.04) 0.00 (0.00) 2.97 (0.14) 0.15 (0.01) HMM-G 7.42 (0.08) 8.99 (0.05) 1.38 (0.05) 2.26 (0.02) 0.79 (0.01) HMM-$\beta$ 14.89 (0.18) 6.16 (0.08) 1.43 (0.03) 2.34 (0.02) 0.81 (0.00) MHMM-G 1.76 (0.04) 4.97 (0.08) 0.94 (0.08) 0.89 (0.07) 0.91 (0.01) MHMM-$\beta$ 0.35 (0.02) 0.63 (0.04) 0.85 (0.11) 3.08 (0.15) 0.45 (0.05) 0.93 (0.01) MHMM-$\beta^{*}$ 500 0.24 (0.01) 0.47 (0.02) 0.00 (0.00) 2.20 (0.10) 0.12 (0.01) HMM-G 7.72 (0.07) 8.19 (0.05) 1.35 (0.03) 2.24 (0.02) 0.82 (0.00) HMM-$\beta$ 14.80 (0.14) 4.15 (0.06) 1.49 (0.03) 2.27 (0.02) 0.81 (0.00) MHMM-G 1.68 (0.03) 4.89 (0.06) 0.55 (0.08) 0.72 (0.07) 0.95 (0.01) MHMM-$\beta$ 0.26 (0.01) 0.50 (0.03) 0.24 (0.07) 2.33 (0.11) 0.36 (0.05) 0.98 (0.01) MHMM-$\beta^{*}$ 1000 0.18 (0.01) 0.35 (0.02) 0.00 (0.00) 1.75 (0.08) 0.13 (0.01) HMM-G 7.68 (0.05) 7.79 (0.03) 1.26 (0.02) 2.21 (0.02) 0.84 (0.00) HMM-$\beta$ 14.08 (0.12) 3.07 (0.05) 1.43 (0.02) 2.22 (0.02) 0.83 (0.00) MHMM-G 1.63 (0.02) 4.81 (0.04) 0.32 (0.08) 0.50 (0.05) 0.97 (0.01) MHMM-$\beta$ 0.20 (0.01) 0.39 (0.02) 0.09 (0.04) 2.05 (0.10) 0.24 (0.04) 0.99 (0.00) MHMM-$\beta^{*}$ 2000 0.13 (0.01) 0.23 (0.01) 0.00 (0.00) 1.21 (0.05) 0.11 (0.01) HMM-G 7.49 (0.04) 7.49 (0.02) 1.10 (0.04) 2.07 (0.04) 0.87 (0.01) HMM-$\beta$ 13.40 (0.08) 2.65 (0.03) 1.31 (0.02) 2.18 (0.02) 0.85 (0.00) MHMM-G 1.62 (0.02) 4.81 (0.03) 0.09 (0.05) 0.29 (0.03) 0.99 (0.00) MHMM-$\beta$ 0.15 (0.01) 0.28 (0.01) 0.05 (0.03) 1.46 (0.07) 0.17 (0.02) 1.00 (0.00)
C.2 Simulation Study on Model Selection
Determining the true number of clusters is crucial. In practice, we advocate the use of both information criteria and domain knowledge. Here, we perform a simulation study to evaluate the performance of different information criteria.
We consider widely used information criteria such as the Akaike Information Criterion (AIC) (Akaike, 1974) and the Bayesian Information Criterion (BIC) (Schwarz, 1978).
$\begin{split}\mbox{AIC}&=-2l(\widehat{\boldsymbol{\Theta}}\mid\boldsymbol{X})+2|\widehat{\boldsymbol{\Theta}}|,\\ \mbox{BIC}&=-2l(\widehat{\boldsymbol{\Theta}}\mid\boldsymbol{X})+\log(NT)|\widehat{\boldsymbol{\Theta}}|.\end{split}$
Here, $N$ represents the total number of QDs, $T$ denotes the length of each QD, $l(\widehat{\boldsymbol{\Theta}}\mid\boldsymbol{X})$ is the observed data log-likelihood, and $|\widehat{\boldsymbol{\Theta}}|$ represents the total number of free parameters in the model.
In addition to AIC and BIC, we also consider the Integrated Completed Likelihood (ICL) criterion (Biernacki et al., 2000), which is commonly used in cluster analysis.
$\begin{split}\mbox{ICL}&=-2l(\widehat{\boldsymbol{\Theta}}\mid\boldsymbol{X},\widehat{\boldsymbol{Z}},\widehat{\boldsymbol{S}})+\log(NT)|\widehat{\boldsymbol{\Theta}}|.\end{split}$
Here, the ICL is an approximation to the complete data likelihood, $l(\widehat{\boldsymbol{\Theta}}\mid\boldsymbol{X},\boldsymbol{Z},\boldsymbol{S})$.
We simulate data under Scenario 3 described above. We then fit the simulated data using MHMM-$\beta$ with varying numbers of clusters and states. For each setting, the model selection is conducted based on AIC, BIC, and ICL.
Table 9 reports the frequency with which each number of clusters is selected by ICL, AIC, and BIC when fitting the MHMM-$\beta$ model over 100 replications. In almost all settings, the correct choice of three clusters dominates, confirming that each criterion can recover the true number of clusters most of the time. Nonetheless, the extent of agreement varies. In balanced scenarios, all three criteria many occasionally over-select the number of clusters, where ICL and BIC favor exactly three clusters more often than AIC. In unbalanced partitions, selecting the correct three-cluster model becomes slightly more challenging; both AIC and BIC more frequently suggest additional clusters, while ICL may also under-select the number of clusters.
These results illustrate the well-known differences in how each criterion penalizes model complexity and cluster separation; while they often converge on three clusters, discrepancies sometimes arise in cases of overlapping or unevenly sized clusters.
Table 9: Proportion (%) of times each number of clusters is selected by ICL, AIC, and BIC under Scenario 3. Rows correspond to different partitions (balanced or unbalanced), sample sizes, and series lengths.
N T ICL AIC BIC
#1 #2 #3 #4 #5 #1 #2 #3 #4 #5 #1 #2 #3 #4 #5
Balanced Partition
50 1000 0 0 81 18 1 0 0 66 33 1 0 0 83 17 0
2000 0 0 74 25 1 0 0 66 29 5 0 0 77 22 1
100 1000 0 0 74 23 3 0 0 69 28 3 0 0 79 21 0
2000 0 0 77 20 3 0 0 74 21 5 0 0 85 15 0
Unbalanced Partition
50 1000 0 12 63 23 2 0 0 59 35 6 0 0 66 29 5
2000 0 0 79 20 1 0 0 71 25 4 0 0 80 20 0
100 1000 0 2 67 24 7 0 0 63 29 8 0 0 68 27 5
2000 0 0 78 18 4 0 0 75 22 3 0 0 85 15 0
Appendix D Additional Results in Data Analysis
D.1 Results from 3-Cluster Model
@@
D.2 Results from 4-Cluster Model
@@
$\begin{split}\hat{\boldsymbol{\pi}}_{1}&=\begin{bmatrix}0.235&0.452&0.313\end{bmatrix},\quad\hat{\boldsymbol{\pi}}_{2}=\begin{bmatrix}0.000&0.125&0.875\end{bmatrix},\\ \hat{\boldsymbol{\pi}}_{3}&=\begin{bmatrix}0.069&0.343&0.588\end{bmatrix},\quad\hat{\boldsymbol{\pi}}_{4}=\begin{bmatrix}0.154&0.573&0.273\end{bmatrix}.\end{split}$
@@
To analyze the relationship between the 3-cluster and 4-cluster solutions, we present the correspondence of cluster assignments as follows.
@@
In the 4-cluster model, Clusters 1 and 2 closely align with Clusters 1 and 2 from the 3-cluster model, with similar estimated transition matrices and equilibrium distributions. Clusters 3 and 4 in the 4-cluster model primarily result from a split of Cluster 3 in the 3-cluster model. Specifically, the new Cluster 3 is associated with more time spent in the high state and exhibits a higher probability of transitioning to and remaining in that state, as indicated by the last column of $\widehat{\boldsymbol{\Pi}}_{3}$. In contrast, Cluster 4 shows longer durations in the medium state and a stronger tendency to remain there compared to Cluster 3.
D.3 Results from 5-Cluster Model
$\widehat{\boldsymbol{\Pi}}_{1}=\begin{bmatrix}0.814&0.160&0.026\\ 0.073&0.774&0.153\\ 0.029&0.280&0.692\end{bmatrix},\quad\widehat{\boldsymbol{\Pi}}_{2}=\begin{bmatrix}0.000&1.000&0.000\\ 0.000&0.993&0.007\\ 0.000&0.000&1.000\end{bmatrix},\quad\widehat{\boldsymbol{\Pi}}_{3}=\begin{bmatrix}0.889&0.092&0.019\\ 0.036&0.785&0.179\\ 0.006&0.132&0.863\end{bmatrix},$
@@ D.3 Results from 5-Cluster Model
$\begin{split}&\hat{\boldsymbol{\pi}}_{1}=\begin{bmatrix}0.237&0.497&0.266\end{bmatrix},\quad\hat{\boldsymbol{\pi}}_{2}=\begin{bmatrix}0.000&0.000&1.000\end{bmatrix},\quad\hat{\boldsymbol{\pi}}_{3}=\begin{bmatrix}0.145&0.364&0.492\end{bmatrix},\\[10.0pt] &\hat{\boldsymbol{\pi}}_{4}=\begin{bmatrix}0.222&0.577&0.201\end{bmatrix},\quad\hat{\boldsymbol{\pi}}_{5}=\begin{bmatrix}0.026&0.425&0.549\end{bmatrix}.\end{split}$
@@
To analyze the relationship between the 4-cluster and 5-cluster solutions, we present the correspondence of cluster assignments as follows.